i need help on these questions i cant find out what i have done wrong i need help on 16 which is 4x+3-2x=2x+7,18 which is 3(x+1)=2x+7
Answer by stanbon(57262)

4x+3-2x=2x+7
2x + 3 = 2x+7
Subtract 2x from both sides to get:
3 = 7
That is a contradiction. There are no
solutions to your equation.
------------------------
3(x+1)=2x+7
3x+1 = 2x+7
x = 6
===========
Cheers,
Stan H.